# Dissipative Yao-Lee Spin-Orbital Model: Exact Solvability and $\mathcal{P}\mathcal{T}$ Symmetry Breaking

### v1

@@ VI Summary and Discussion
Furthermore, we break away from sectors that host NESS and analytically compute the Liouvillian eigenvalues in a translationally-invariant symmetry sector. We discover a three-stage behavior of the eigenvalues as the dissipation strength $\gamma$ is increased: first, below a system-size dependent threshold $\gamma_{\text{PT}}$, all eigenvalues are imaginary, corresponding to a $\mathcal{P}\mathcal{T}$ preserving phase. For $\gamma_{\text{PT}}<\gamma<\gamma^{*}$, the spectrum contains both purely real and purely imaginary eigenvalues, and the system is in a $\mathcal{P}\mathcal{T}$-mixed phase [quantum_quench_drivendissipative]. Finally, when the dissipation strength is larger than $\gamma^{*}$, all modes have real eigenvalues and decay with distinct rates, i.e. the system is $\mathcal{P}\mathcal{T}$ broken. We are able to obtain a qualitative (from oscillatory to decaying) crossover of observable dynamics, up to a decaying envelope.
We supplement our analysis with numerical calculations in both real and momentum space. In $k$-space, we discover that the exceptional points form an exceptional ring [exceptional_spin_liquid]. In real space, we calculate the Liouvillian eigenvalues and demonstrate that the scalings of $\gamma_{\text{PT}}$ and $\gamma^{*}$ with system size are in agreement with our analysis.
There are a few interesting directions for future exploration. First, in the long-time limit, the relaxation timescale is determined by the Liouvillian gap, which is the slowest non-zero decay rate and governs relaxation dynamics at long times in the system [liouvillian_gap_def]. More precisely, since every eigenvalue $\lambda_{n}$ of $\mathcal{L}$ has a non-positive real part [third_quantization], we can order the Liouvillian spectrum by its real part as $0=|\Re\lambda_{0}|<|\Re\lambda_{1}|<\cdots<|\Re\lambda_{n}|$. The Liouvillian gap is defined as [liouvillian_gap_def]
@@
One can analyze “excitations” on top of the sectors that host steady states: in particular, the model hosts distinct excitations in the Liouvillian spectrum, corresponding to flipping a vertical plaquette $V_{p}$, flipping a hexagonal plaquette $W_{p}$, or creating/removing fermions. Note that since the fermion number, instead of parity [Shackleton_dissipative, arovas_dissipative], is conserved, there is an extensive number of distinct excitations in the fermionic degree of freedom.
@@
From Sec. IV.2, when we analyzed the nature of steady states for $\gamma=0$ and $\gamma=\infty$, we have already seen a hint of this. Without dissipation, the system’s steady states are pure states; at very large dissipation, $\rho_{\text{ss}}$ are fully mixed (i.e. identity) in the $\sigma$ sector. Therefore, there might be a purity transition of the steady states as the dissipation strength increases [Huber_PT].
Finally, it is known from non-Hermitian physics that exceptional points in two dimensions and above have distinct topology compared to those in 1D [nonhermitian_bandtheory]. While exceptional points in the Liouvillian spectrum have been studied in one-dimensional spin and fermionic systems, their nature has not been elucidated in a two-dimensional, dissipative setting. Our exactly solvable dissipative model, therefore, provides a playground for exploring the topology of Liouvillian exceptional points in a higher dimension.
@@ Appendix A Conditions for $\mathcal{P}\mathcal{T}$ Symmetric Liouvillian
On the super-operator level, a Liouvillian $\mathcal{L}$ is $\mathcal{P}\mathcal{T}$ symmetric if it satisfies [prosen_PT]:
@@
The above definition is generic on the super-operator level. On the operator level, it turns out that $\mathcal{L}$ is $\mathcal{P}\mathcal{T}$ symmetric if the Hamiltonian $H$ and jump operators $\{L_{i}\}$ satisfy the following conditions [prosen_PT_criteria]:
@@
Our dissipative spin-orbital model satisfies all of the above conditions. The first condition is immediately met, as $\{\sigma_{i}^{x},(\sigma_{i}^{x})^{\dagger}\}=\{\sigma_{i}^{z},(\sigma_{i}^{z})^{\dagger}\}=2\mathbb{I}$ for all sites $i$. Under conjugation by the unitary $U$ defined in Eq. (36), the Hamiltonian (Eq. (6)) (which is real, $H=H^{*}$) remains invariant. While conjugation by $U$ flips the sign of individual $\sigma_{i}^{x}$ and $\sigma_{i}^{z}$, $H$ only contains two-body terms $\sigma_{i}^{x}\sigma_{j}^{x}$ and $\sigma_{i}^{z}\sigma_{j}^{z}$ in the spin channel, and therefore condition (2) is satisfied. Finally, since $U\sigma_{i}^{z}U^{\dagger}=-\sigma_{i}^{z}$, $U\sigma_{i}^{x}U^{\dagger}=-\sigma_{i}^{x}$, and the jump operators are Hermitian, $L_{i}=L_{i}^{\dagger}$, condition (3) is satisfied, with $V=\mathbb{I}$. Therefore the Liouvillian in our model is indeed $\mathcal{P}\mathcal{T}$ symmetric.

### v2

@@ VI Summary and Discussion
Furthermore, we break away from sectors that host NESS and analytically compute the Liouvillian eigenvalues in a translationally-invariant symmetry sector. We discover a three-stage behavior of the eigenvalues as the dissipation strength $\gamma$ is increased: first, below a system-size dependent threshold $\gamma_{\text{PT}}$, all eigenvalues are imaginary, corresponding to a $\mathcal{P}\mathcal{T}$ preserving phase. For $\gamma_{\text{PT}}<\gamma<\gamma^{*}$, the spectrum contains both purely real and purely imaginary eigenvalues, and the system is in a $\mathcal{P}\mathcal{T}$-mixed phase [78]. Finally, when the dissipation strength is larger than $\gamma^{*}$, all modes have real eigenvalues and decay with distinct rates, i.e. the system is $\mathcal{P}\mathcal{T}$ broken. We are able to obtain a qualitative (from oscillatory to decaying) crossover of observable dynamics, up to a decaying envelope.
We supplement our analysis with numerical calculations in both real and momentum space. In $k$-space, we discover that the exceptional points form an exceptional ring [85]. In real space, we calculate the Liouvillian eigenvalues and demonstrate that the scalings of $\gamma_{\text{PT}}$ and $\gamma^{*}$ with system size are in agreement with our analysis.
There are a few interesting directions for future exploration. First, in the long-time limit, the relaxation timescale is determined by the Liouvillian gap, which is the slowest non-zero decay rate and governs relaxation dynamics at long times in the system [52]. More precisely, since every eigenvalue $\lambda_{n}$ of $\mathcal{L}$ has a non-positive real part [61], we can order the Liouvillian spectrum by its real part as $0=|\Re\lambda_{0}|<|\Re\lambda_{1}|<\cdots<|\Re\lambda_{n}|$. The Liouvillian gap is defined as [52]
@@
One can analyze “excitations” on top of the sectors that host steady states: in particular, the model hosts distinct excitations in the Liouvillian spectrum, corresponding to flipping a vertical plaquette $V_{p}$, flipping a hexagonal plaquette $W_{p}$, or creating/removing fermions. Note that since the fermion number, instead of parity [73, 29], is conserved, there is an extensive number of distinct excitations in the fermionic degree of freedom.
@@
From Sec. IV.2, when we analyzed the nature of steady states for $\gamma=0$ and $\gamma=\infty$, we have already seen a hint of this. Without dissipation, the system’s steady states are pure states; at very large dissipation, $\rho_{\text{ss}}$ are fully mixed (i.e. identity) in the $\sigma$ sector. Therefore, there might be a purity transition of the steady states as the dissipation strength increases [32].
Finally, it is known from non-Hermitian physics that exceptional points in two dimensions and above have distinct topology compared to those in 1D [74]. While exceptional points in the Liouvillian spectrum have been studied in one-dimensional spin and fermionic systems, their nature has not been elucidated in a two-dimensional, dissipative setting. Our exactly solvable dissipative model, therefore, provides a playground for exploring the topology of Liouvillian exceptional points in a higher dimension.
@@ Appendix A Conditions for $\mathcal{P}\mathcal{T}$ Symmetric Liouvillian
On the super-operator level, a Liouvillian $\mathcal{L}$ is $\mathcal{P}\mathcal{T}$ symmetric if it satisfies [63]:
@@
The above definition is generic on the super-operator level. On the operator level, it turns out that $\mathcal{L}$ is $\mathcal{P}\mathcal{T}$ symmetric if the Hamiltonian $H$ and jump operators $\{L_{i}\}$ satisfy the following conditions [62]:
@@
Our dissipative spin-orbital model satisfies all of the above conditions. The first condition is immediately met, as $\{\sigma_{i}^{x},(\sigma_{i}^{x})^{\dagger}\}=\{\sigma_{i}^{z},(\sigma_{i}^{z})^{\dagger}\}=2\mathbb{I}$ for all sites ii. Under conjugation by the unitary $U$ defined in Eq. (36), the Hamiltonian (Eq. (6)) (which is real, $H=H^{*}$) remains invariant. While conjugation by $U$ flips the sign of individual $\sigma_{i}^{x}$ and $\sigma_{i}^{z}$, $H$ only contains two-body terms $\sigma_{i}^{x}\sigma_{j}^{x}$ and $\sigma_{i}^{z}\sigma_{j}^{z}$ in the spin channel, and therefore condition (2) is satisfied. Finally, since $U\sigma_{i}^{z}U^{\dagger}=-\sigma_{i}^{z}$, $U\sigma_{i}^{x}U^{\dagger}=-\sigma_{i}^{x}$, and the jump operators are Hermitian, $L_{i}=L_{i}^{\dagger}$, condition (3) is satisfied, with $V=\mathbb{I}$. Therefore the Liouvillian in our model is indeed $\mathcal{P}\mathcal{T}$ symmetric.